### v2

# Multiscale Growth Kinetics of Model Biomolecular Condensates Under Passive and Active Conditions

Abstract
Living cells exhibit a complex organization comprising numerous compartments, among which are RNA- and protein-rich membraneless, liquid-like organelles known as biomolecular condensates. Energy-consuming processes regulate their formation and dissolution, with (de-)phosphorylation by specific enzymes being among the most commonly involved reactions. By employing a model system consisting of a phosphorylatable peptide and homopolymeric RNA, we elucidate how enzymatic activity modulates the growth kinetics and alters the local structure of biomolecular condensates. Under passive condition, time-resolved ultra-small-angle X-ray scattering with synchrotron source reveals a nucleation-driven coalescence mechanism maintained over four decades in time, similar to the coarsening of simple binary fluid mixtures. Coarse-grained molecular dynamics simulations show that peptide-decorated RNA chains assembled shortly after mixing constitute the relevant subunits. In contrast, actively-formed condensates initially display a local mass fractal structure, which gradually matures upon enzymatic activity before condensates undergo coalescence. Both types of condensate eventually reach a steady state but fluorescence recovery after photobleaching indicates a peptide diffusivity twice higher in actively-formed condensates consistent with their loosely-packed local structure. We expect multiscale, integrative approaches implemented with model systems to link effectively the functional properties of membraneless organelles to their formation and dissolution kinetics as regulated by cellular active processes.
[…]
{tocentry}
1 Introduction
Membraneless organelles are intracellular compartments devoid of lipid membrane, consisting of multiple proteins and possibly nucleic acids (DNA and/or RNA), and formed by liquid-liquid phase separation 1, 2, 3, 4. These biomolecular condensates perform essential biological functions by sequestering specific factors and promoting biochemical reactions following the needs of the cell 5. They also serve as protocells for the study of the emergence of functional cellular life or as catalytic microcompartments 6, 7. Many condensates found in eucaryotic cells contain RNA and are involved in gene expression, such as nucleoli 8, P granules 9, stress granules 10 or Cajal bodies 11 to name a few. Protein-RNA condensates may also play a role in pathological processes 12, for example by facilitating the transcription of viral RNA during SARS-Cov-2 replication 13. Membraneless organelles are relatively short-lived, in that they are formed to perform specific functions and are dissolved after the completion of these tasks. Such a flexibility is maintained through energy-consuming enzymatic processes. In particular, various post-translational modifications have been shown to regulate intracellular phase separation, including phosphorylation 14, methylation 15, acetylation and combinations thereof 16.
Phase separation is a phenomenon by which a homogeneous material, spontaneously or not, separates into two or more distinct domains (phases) with different physical and/or chemical properties. Extensive experimental and theoretical efforts have been devoted to elucidating the phase-separation kinetics of binary systems 17, including fluid mixtures relevant to biomolecular condensates, under both passive (spontaneous) and active (chemically driven) conditions 18, 19, 20. A major limitation in these investigations lies in the lack of information across length and time scales spanning several orders of magnitude. Indeed, biomolecular condensates are mostly micrometer-sized and therefore, fluorescence microscopy has become the standard technique to monitor their life cycle, especially within cells 8. Quite recently, holographic microscopy has demonstrated its ability to monitor the growth of PopZ-cation condensates without the need for protein labeling 21. In all these cases, submicrometer-sized condensates formed along the growth pathway are overlooked, and events occurring on subsecond timescales often remain inaccessible. Conversely, coarse-grained molecular dynamics simulations can provide a wealth of information at the molecular level 22, and a few phase-separated protein-RNA condensates have been elucidated 23, 24. Nonetheless, coarse-grained models face substantial challenges in quantitatively reproducing the thermodynamic properties of biomolecular condensates. In addition, although they allow the simulation of relatively large supramolecular systems, the accessible timescales remain limited to a few microseconds at most.
To bridge these multiple scales, time-resolved small-angle X-ray scattering (TR-SAXS) using synchrotron source has emerged as a state-of-the-art technique, enabling to track self-assembly processes with millisecond temporal resolution and spatial sensitivity ranging from a few to several hundred nanometers 25, 26, 27, 28, 29. Within the context of liquid-liquid phase separation, the assembly kinetics of a prion-like domain has revealed the existence of a double nucleation barrier to overcome before the condensate growth can proceed 30. With the advent of time-resolved ultra-small-angle X-ray scattering (TR-USAXS) 31 available at a very few synchrotron facilities worldwide, the capabilities of X-ray scattering have extended to the micrometer scale. For example, the spinodal decomposition of bovine serum albumin induced by YCl3 was probed during the first two minutes following the mixing of components 32. Likewise, coacervates composed of oppositely-charged polyelectrolytes were found to change from mass fractal to surface fractal morphologies during their growth driven by coalescence 33. Nevertheless, how enzymatic activity regulates the growth mechanism, the local structure, and the morphology of biomolecular condensates – compared to conditions lacking active processes – remains experimentally unaddressed.
To fill this knowledge gap, we employ a model system of biomolecular condensates whose formation is triggered by the dephosphorylation of serine residues, thereby mimicking the enzymatic regulation encountered in intracellular membraneless organelles 34. This system consists of an arginine-rich peptide (RRASLRRASL) containing phosphorylatable serines, associated with poly(uridylic acid) (polyU) RNA. Condensates form spontaneously when the serines bear positive charges – a condition hereafter referred to as passive – or are enzymatically induced by lambda phosphatase (LPP), which removes phosphate groups from initially phosphorylated serines, defining the active condition. This flexible system enables the comparison between condensate formation mechanisms under passive and active conditions. We therefore combine TR-USAXS experiments performed at varying peptide and enzyme concentrations, with confocal microscopy and coarse-grained molecular dynamics simulations to capture these mechanisms across a wide range of length- and timescales.
2 Results
2.1 Phase separation occurs in the cationic regime
Upon mixing RRASLRRASL peptide with anionic polyU RNA (Fig. 1a) in 4 mM MgCl2, 2.25 mM ATP, 1.6 mM dithiothreitol, 50 mM HEPES pH 7.4 at 37 °C, biomolecular condensates in the form of liquid-like droplets are spontaneously made by coacervation 34. A direct evidence for the presence of micrometer-sized droplets is given by the turbid aspect of the solution (Fig. 1b). Confocal microscopy images performed with 0.5% of fluorescent TAMRA-peptide confirm the spherical shape and the micrometer-scale size of the condensates (Fig. 1c). At a given RNA concentration, turbidity measurements display a clear peptide threshold (saturation) concentration (Fig. 1d), which is characteristic of phase separation. It is worth noting that the phase boundary is almost superimposed with the electroneutrality line. Phase separation occurs in the cationic regime, when the total positive charges supplied by the peptides outnumber the total negative charges carried by RNA.
Figure 1: (a) Schematic representation of RRASLRRASL peptide (red) and polyU RNA (blue) forming condensates upon Coulomb interaction. (b) A solution of condensates in a quartz cuvette exhibits a high level of turbidity. (c) Confocal microscopy image of condensates obtained with 1,000 µM of peptide including 0.5% of TAMRA-peptide, and 0.5 g.L-1 of RNA. (d) Contour map of turbidity measured at 500 nm as a function of peptide and RNA concentrations. The red dashed line indicates electroneutrality.
The SAXS measurement of peptides reveals no sign of aggregation (Fig. LABEL:fig:SI-saxs_peptide) given that the scattering curve at low $q$-values – $q$ being the magnitude of the scattering wavevector – is flat, which means that the Guinier plateau is reached. Furthermore, the measured radius of gyration ($R_{\mathrm{g}}=\text{0.9\ nm}$) is consistent with the small number of amino acids (ten). Likewise, the scattering curve of phosphorylated peptides reveals a radius of gyration of 0.84 nm (Fig. LABEL:fig:SI-saxs_peptide). PolyU RNA has $250\pm 80$ (s.d.) nucleotides according to mass photometry (Fig. LABEL:fig:SI-mass_photometry_polyU) and its radius of gyration is estimated to be 11.6 nm (Fig. LABEL:fig:SI-saxs_polyU). The Kratky representation of the RNA scattering curve displays an increase of the normalized intensity $(qR_{\mathrm{g}})^{2}(I/I_{0})$ at high $qR_{\mathrm{g}}$ – where $I$ stands for the scattering intensity and $I_{0}$ is the forward scattering intensity –, which indicates a stretched conformation as expected for a fully charged polyelectrolyte.
2.2 Passively-formed condensates grow by nucleated Brownian motion-induced coalescence
Due to the Coulomb interactions underlying coacervation, growth is expected to start as early as a few milliseconds following the mixing of components, spanning lengthscales from several tens of nanometers to a few micrometers. Accordingly, TR-USAXS experiments on synchrotron source are carried out with a very large sample-to-detector distance of 31 m, and the rapid mixing is performed with a stopped-flow device that enables a dead time as short as 2.5 ms. Figure 2a depicts scattering curves measured at different times for a peptide concentration of 1,000 µM (see Fig. LABEL:fig:SI-tr-saxs_passive for other peptide concentrations). The coacervation begins immediately after mixing and the remarkably flat Guinier plateau at low $q$-values is characteristic of the absence of interaction between condensates. Note that these measurements are repeatable from the first milliseconds up to nearly 10 min (Fig. LABEL:fig:SI-comp_growth_kinetics_1000uM).
Figure 2: TR-USAXS measurements under passive conditions. (a) Scattering curves at different times of a solution containing 1,000 µM of peptide and 0.5 g.L-1 of RNA. (b) Mean radius $\langle R\rangle$ of condensates as a function of time for various peptide concentrations. The numbers in brackets are the peptide-to-RNA charge ratios. $\langle R\rangle$ is estimated by fitting the scattering curves with a model of lognormally-distributed spheres. The dashed line represents a scaling law in $t^{1/3}$. (c) Arbitrarily scaled number density of condensates versus time at various peptide concentrations. The dashed line is a scaling law in $t^{-1}$. The RNA concentration in (b) and (c) is 0.5 g.L-1 except for the highest peptide concentration of 2,000 µM where the RNA concentration is 1 g.L-1.
Since confocal microscopy clearly indicates a spherical morphology for all the condensates (Fig. 1c), we fit each scattering curve with a model of polydisperse spheres expressed as:
$I(t,q)=\frac{I_{0}(t)}{\langle R^{6}(t)\rangle}\displaystyle\int_{0}^{+\infty}\left[3\frac{\sin(qR)-qR\cos(qR)}{(qR)^{3}}\right]^{2}p(R;\langle R(t)\rangle,\Delta R(t))R^{6}dR+C_{\mathrm{bkg}}$ (1)
where $R$ is the condensate radius and $p(R;\langle R(t)\rangle,\Delta R(t))$ the size distribution assumed here to be lognormal with mean radius $\langle R(t)\rangle$ and width $\Delta R(t)$. $C_{\mathrm{bkg}}$ is a constant accounting for the background. The three fitting parameters at each time step are therefore $I_{0}$, $\langle R\rangle$ and $\Delta R$, $\langle R^{6}\rangle$ being computed with the size distribution. Figure 2b shows the time evolution of $\langle R\rangle$ inferred from the scattering curves at various peptide concentrations. Strikingly, the mean radius obeys a scaling law of the form $\langle R\rangle=(Kt)^{1/3}$ – where $K$ is a dynamic prefactor –, as early as the first few tens of milliseconds, and this behavior persists over nearly four decades in time for the highest peptide concentrations. An exponent of $1/3$ during liquid-liquid phase separation is the hallmark of either17, 19: (i) diffusion-limited Ostwald ripening where condensate-forming molecules migrate from small condensates to large condensates; or (ii) Brownian motion-induced coalescence during which condensates fuse to form increasingly larger condensates. In the latter case, the dynamic prefactor is expressed as $K\simeq 6k_{\mathrm{B}}T\phi/5\pi\eta$, with $k_{\mathrm{B}}$ the Botzmann constant, $T$ the temperature, $\phi$ the condensate volume fraction, and $\eta$ the solvent viscosity19. Whereas $K$ depends on the condensate volume fraction for the coalescence mechanism, it is only related to the properties of condensate-forming molecules such as the molar volume and the saturation concentration in the case of Ostwald ripening. The ratio of the measured dynamic prefactors for experiments performed with peptide concentrations of 1,000 and 2,000 µM while maintaining the peptide-to-RNA charge ratio constant, is estimated to be 1.6. Although a factor 2 is theoretically expected, the fact that $K$ increases with volume fraction suggests that coalescence mostly drives the growth kinetics.
A quantity proportional to the number density of condensates $N$ can be deduced from the forward scattering intensity $I_{0}\equiv I(q\rightarrow 0)$ and the size distribution at each time since $I_{0}\sim N\langle V^{2}\rangle\sim N\langle R^{6}\rangle$ with $V$ the condensate volume 35. Figure 2c depicts this quantity versus time and demonstrates that $N$ scales as $t^{-1}$, collapsing onto a single curve at late times irrespective of the peptide concentration. This behavior is consistent with a coalescence mechanism: the volume fraction $\phi$ remains constant over time, and applying the scaling law for coalescence, $R^{3}\sim\phi t\sim NR^{3}t$, leads to $N\sim t^{-1}$, with a prefactor that is indeed independent of $\phi$.
Additionally, Figure 2b illustrates the nucleation preceding the condensate growth. As the peptide concentration approaches the saturation concentration ($\approx$400 µM), an increasingly long induction time is observed before a sufficient signal-to-noise ratio on the scattering intensity enables to infer condensate size (see the 400 and 600 µM curves on Fig. 2b). A close examination of the SAXS curves at 400 µM reveals no detectable evolution until $t=32\text{ s}$, while within the following seconds, the scattering intensity dramatically increases. We hypothesize that prior to nucleation, each isolated RNA chain is decorated with bound peptides. This interpretation is supported by a comparison between the average scattering intensity measured prior to nucleation and the sum of the individual contributions from peptides and RNA. As the former exceeds the latter (Fig. LABEL:fig:SI-supersaturation), peptides are indeed bound to RNA. From the forward scattering intensity, we estimate that approximately 30 peptides are associated with each RNA chain (see Supporting Information). Furthermore, fitting the average scattering intensity using a Fisher–Burford model (Fig. LABEL:fig:SI-supersaturation) 36, which describes simple fractal aggregates, yields a mass fractal dimension of 1.4. This value indicates that the clusters adopt an extended, chain-like conformation, with a size ($R_{\mathrm{g}}=13\ \text{nm}$) comparable to that of bare RNA ($R_{\mathrm{g}}=11.6\ \text{nm}$).
To get deeper insights into the early events at the molecular level, we perform coarse-grained molecular dynamics (MD) simulations with implicit solvent by using a recently released CALVADOS RNA force field 37. The numerical model is able to reproduce a saturation concentration of 350 µM (Fig. LABEL:fig:SI-computed_phase_diagram), that is, within the range of values found experimentally (Fig. 1d; $\simeq$400 µM). Peptides and RNA are initially mixed in a simulation box, and Figure 3 depicts the time evolution of the number of peptides contained in the largest peptide-RNA cluster. Very rapidly, all the peptides bind on RNA, which are thus decorated with an average of 50 peptides. Remarkably, the total number of bound peptides barely varies afterwards (Fig.LABEL:fig:SI-MD_largest_cluster_allc), confirming the role of peptide-decorated RNA as stable subunits. The peptide-RNA clusters further grow in a stepwise manner by coalescence. This is illustrated on Fig.3 wherein the steps observed in the number of peptides in the largest cluster become higher over the course of the simulation. In particular, snapshots 4 and 5 show two clusters comprising 400 and 200 peptides, respectively, that is, 8 and 4 RNA chains (Fig. LABEL:fig:SI-MD_largest_cluster_allb), respectively, merging into a single cluster. Note that fusion events are increasingly rare with the simulation time steps as the diffusion of larger clusters is much slower.
Figure 3: Coarse-grained MD simulation of passively-formed condensates. The graph plots the number of peptides in the largest cluster as a function of the simulation time steps in logarithmic scale. The snapshots illustrate the coalescence events as labeled on the graph. Peptides (800 µM) are in green and RNA (0.5 g.L-1) in purple.
At a macroscopic level, condensates exceeding several micrometers in size cannot be probed even by USAXS. Therefore, we carry out confocal microscopy imaging by using again 0.5% of fluorescent TAMRA-peptide. Figure 4a depicts a low magnification view of condensates at a peptide concentration of 1,000 µM, 25 min after mixing with RNA. The radius distribution, shown in Fig. 4b, resembles a lognormal distribution, as expected from a coalescence mechanism 19. No macroscopic phase separation is observed, most likely as a result of Coulomb repulsion that self-regulates the growth of condensates 21.
Figure 4: (a) Confocal microscopy image of condensates obtained with a peptide concentration of 1,000 µM including 0.5% of TAMRA-peptide and 0.5 g.L-1 of RNA. The image is acquired 25 min after mixing components at 37 °C. (b) Condensate radius distribution ($N=707$) inferred from a larger view of (a). The red line is a fit with a lognormal distribution.
2.3 Network-like local structure under active condition
In the subsequent series of experiments, we examine the growth kinetics of condensates under active condition. The serine residues of the RRASLRRASL peptide are initially phosphorylated, which results in a neutral net charge. As a consequence, mixing phosphorylated peptides and RNA does not give rise to phase separation. Phase separation is instead triggered by the action of lambda phosphatase (LPP), which cleaves the phosphate groups from phosphorylated peptides. Figure 5a plots scattering curves collected under the same concentrations as in Fig. 2a in the presence of 800 U.mL-1 LPP and using initially double phosphorylated peptide (see Fig. LABEL:fig:SI-tr-saxs-active for other peptide concentrations). The condensate formation is much slower than under passive condition because phase separation is now limited by the LPP enzymatic activity. Secondly, the scattering curves in the intermediate $q$-values, i.e., between 10-2 and 10-1 nm-1, differ markedly. The inset of Figure 5a presents a comparison between two scattering curves with similar $I_{0}$ values. For actively-formed condensates, the decay of $I(q)$ at small and medium $q$-values is moderate ($I\sim q^{-2.4}$), which suggests a local mass fractal structure akin to that found in gels, while for passively-formed condensates, the decay slope is more consistent with a surface fractal structure. Therefore, instead of fitting the scattering curves with a model of polydisperse spheres as earlier, we apply a model of polydisperse mass fractal clusters with a compressed exponential cutoff function 38, 36. The form factor of the clusters is expressed by:
$F(q;\xi)=\displaystyle\int_{0}^{+\infty}r^{D-1}h(r;\xi)\frac{\sin(qr)}{qr}dr\big/\int_{0}^{+\infty}r^{D-1}h(r;\xi)dr$ (2)
where $D$ stands for the mass fractal dimension, and $h(r;\xi)$ is a cutoff function that describes the perimeter of the cluster of size $\xi$ and which must decay faster than any power law. The denominator ensures the normalization condition, namely, $F(q\rightarrow 0;\xi)=1$. We choose a generic exponential form for the cutoff function 36, i.e., $h(r;\xi)=e^{-(r/\xi)^{\beta}}$ with $\beta=3$ as it fits satisfactorily the scattering curves at late times. As before (Eq. 1), the scattering intensity is computed as a convolution product of the form factor and a lognormal size distribution defined by $p(\xi;\langle\xi(t)\rangle,\Delta\xi(t))$, so that
$I(t,q)=\frac{I_{0}(t)}{\langle\xi^{2D(t)}(t)\rangle}\displaystyle\int_{0}^{+\infty}F(q;\xi)p(\xi;\langle\xi(t)\rangle,\Delta\xi(t))\xi^{2D(t)}d\xi+C_{\mathrm{bkg}}$ (3)
In practice, the scattering curves are fitted over the region $q<\text{0.1 nm\textsuperscript{-1}}$, which corresponds to lengthscales above the size of a single RNA chain. The blue dashed lines of Fig. 5a represents fits to the experimental data (black solid lines), providing estimates of $I_{0}(t)$, $\langle\xi(t)\rangle$, $\Delta\xi(t)$ and $D(t)$ for each condition. Figure 5b depicts the time evolution of $\langle\xi\rangle$ for three peptide concentrations. We observe mostly two steps: (i) Condensates with size varying around 100 nm are formed from the first minutes after mixing. These condensates display a network-like local structure as evidenced by a mass fractal dimension $D$ well below 3 (Fig. 5c). After some period of maturation during which $D$ increases steadily while $\langle\xi\rangle$ remains unchanged, the condensates abruptly start to increase in size, with $\langle\xi\rangle$ seemingly converging toward the classical coarsening behavior scaling as $t^{1/3}$ (black dashed line in Fig. 5b). During the latter stage, $D$ remains stable above 2.5, which indicates that the condensates are fairly compact. Notice that at 600 µM peptide concentration, no growth is observed for at least the first 2,000 s. Increasing the LPP concentration from 800 U.mL-1 to 1,200 U.mL-1 (gray discs in Figs. 5b and c) shortens the maturation stage consistently with higher enzymatic activity. Finally, it is worth mentioning that although the condensate polydispersity $\Delta\xi/\langle\xi\rangle$ varies significantly during the maturation stage, it tends to a nearly constant value during the growth stage (Fig. LABEL:fig:SI-polydispersity_active), consistently with the self-similar size distribution encountered classically in coarsening mechanisms.
Figure 5: TR-USAXS measurements under active conditions. (a) Scattering curves at various times (black solid lines) fitted with a polydisperse mass fractal model (blue dashed lines). The solution contains 1,000 µM of initially phosphorylated peptide, 0.5 g.L-1 of RNA and 800 U.mL-1 of LPP. The inset compares two scattering curves collected with peptide and RNA concentrations of 1,000 µM and 0.5 g.L-1, respectively, under passive (gray dashed line) and active (black solid line; 800 U.mL-1 of LPP), at times where the $I_{0}$ values are similar. (b) Mean condensate size $\langle\xi\rangle$ as a function of time for various peptide concentrations. LPP concentration is 800 U.mL-1 except in one experiment (gray discs; 1,200 U.mL-1). The dashed line indicates a scaling law in $t^{1/3}$. (c) Mass fractal dimension $D$ versus time for the same experimental conditions as in (b).
MD simulations that explicitly account for enzymatic activity on phosphorylated peptides are not computationally feasible due to the inherently slow timescales of the reactions involved. Instead, we model cluster formation in systems containing distinct fixed peptide compositions, thereby mimicking different stages of phosphatase-driven dephosphorylation. Figure 6a presents the composition of the six simulated systems, ranging from S1, which predominantly contains double phosphorylated peptides, to WT, which consists solely of non-phosphorylated peptides as set under passive conditions. The exact peptide compositions are detailed in Table LABEL:tab:SI-composition_condensate_active. As shown in Fig. 6b, decreasing the phosphorylation level enhances the ability of clusters to incorporate RNA chains within the same simulation timeframe. This behavior is expected as coacervation relies on positively charged peptides to compensate for the negative charges of RNA. More intriguingly, the final clusters display increasingly loose and heterogeneous morphologies when a substantial fraction of double or mono phosphorylated peptides is present, in contrast to the compact aggregates formed by non-phosphorylated peptides (Fig. 6c). Figures 6d and e quantitatively support this observation by showing that both the radius of gyration and the asphericity of the final clusters increase with peptide phosphorylation level, regardless of the number of incorporated RNA chains. In summary, higher phosphorylation levels lead to less compact and more extended cluster structures, in agreement with experimental measurements.
Figure 6: Coarse-grained MD simulations of condensate growth mimicking active conditions. (a) Peptide compositions of the six simulated systems. (b) Time evolution of the number of RNA chains incorporated into the largest cluster for each system. Shaded areas represent standard deviations over three simulations. (c) Representative snapshots of clusters for the WT and S3 systems. (d) Radius of gyration ($R_{gyr}$) of clusters containing one, two, or three RNA chains for different peptide compositions. (e) Asphericity of clusters containing one, two, or three RNA chains for different peptide compositions.
Next, we investigate the evolution of actively-formed condensates over extended length- and timescales by confocal microscopy. Figures 7a-d depicts confocal microscopy images acquired at various times for a peptide concentration of 1,000 µM. Figure 7e represents the time evolution of the mean radius $\langle R\rangle$ and the standard deviation calculated from image analysis. $\langle R\rangle$ increases until 60 min and remains constant afterwards. Once again, the condensates attain a steady state due to self-regulation driven by Coulomb repulsion. Note that the radii deduced by confocal microscopy (Fig. 7e) are underestimated compared to those obtained by X-ray scattering (Fig. 5b) because the sizes are near the resolution limit of the confocal setup. At a lower peptide concentration of 600 µM, confocal microscopy reveals that 2 h are necessary for the condensate mean radius to reach 1 µm (Fig. LABEL:fig:SI-confocal_active_600_uM). Interestingly, the normalized size distributions, $R/\langle R\rangle$, computed at various times, are mostly superimposed, which supports the self-similarity hypothesized earlier with TR-USAXS data. Additionally, these distributions are slightly asymmetric, resembling the lognormal distribution expected for a coalescence mechanism 19. This is further illustrated by Figs. 7g and h showing three coalescence events occurring over 4 min.
Figure 7: Confocal microscopy of actively-formed condensates. Initial phosphorylated peptide concentration is 1,000 µM including 0.5% of TAMRA-peptide, and LPP concentration is 800 U.mL-1. (a-d) Condensate images acquired at various times after mixing components. (e) Mean condensate radius $\langle R\rangle$ versus time obtained from image analysis. Error bars are standard deviations. (f) Normalized size distribution $R/\langle R\rangle$ at various times during growth and steady-state stages. Black dashed line is a lognormal distribution plotted for comparison. (g-h) Coalescence events between condensates observed at three different locations (white boxes).
2.4 Actively-formed condensates exhibit faster nanoscale dynamics
Intrigued by the difference of local structures uncovered by X-ray scattering, we carry out fluorescence recovery after photobleaching (FRAP) measurements to estimate the diffusivity of TAMRA-peptide within both passively- and actively-formed condensates in steady state. Figures 8a-d demonstrate the liquid-like nature of condensates with a fluorescence recovery obtained within just a few seconds. By integrating the fluorescence intensity over condensates and fitting the time evolution with an exponential decay function, the diffusivity of TAMRA-peptide, $D_{\mathrm{c}}$, is estimated (Figs. 8e and f). TAMRA-peptide diffuses twice as fast within actively-formed condensates (0.024 µm2.s-1) as within passively-formed condensates (0.012 µm2.s-1). This finding agrees well with the loosely-packed, network-like local structure measured for actively-formed condensates, which confers a facilitated translational mobility to peptides.
Figure 8: FRAP measurements on passively- and actively-formed condensates in steady state. Peptide concentration is 1,000 µM including 0.5% of TAMRA-peptide. For active condition, 800 U.mL-1 of LPP is used. (a-d) Confocal microscopy images of an actively-formed condensate at various times after photobleaching. Normalized fluorescence intensity versus time for passively- (e) and actively-formed (f) condensates. A dozen condensates are used for each condition. Error bars are standard deviations. The diffusivities $D_{\mathrm{c}}$ are deduced by fitting the data with an exponential decay function (red line).
3 Conclusion
The kinetics of the liquid–liquid phase separation leading to the formation of biomolecular condensates composed of a cationic, phosphorylatable peptide and polyU RNA – under both passive and active conditions – are compared using multiscale approaches that span up to four decades in time and bridge spatial scales from the molecular to the micrometer level. Based on the results, Figure 9 proposes two growth mechanisms depending on whether enzymatic activity is involved or not.
Figure 9: Schematics of the growth mechanisms for passively- and actively-formed biomolecular condensates.
For passively-formed condensates, peptides bear a positive net charge and bind on RNA in less than a millisecond after mixing. Peptide-decorated RNA then serve as robust subunits and behave as an immiscible fluid in a solvent. After nucleation, spherical condensates steadily grow following the classical coarsening scaling law in $\langle R\rangle=(Kt)^{1/3}$. The prefactor $K$, as well as confocal microscopy images and MD simulations suggest a Brownian motion-induced coalescence mechanism. Eventually, coalescence is impaired by Coulomb repulsion, leading to steady state. Coalescence is preferred over ripening likely because transferring large peptide-decorated RNA from small to large condensates would be too costly in terms of free energy.
As for actively-formed condensates, double phosphorylated peptides bear a zero net charge and bind on RNA upon phosphate cleavage by LPP. Resultantly, RNA chains are slowly crosslinked by dephosphorylated peptides making up loosely-packed condensates with a network-like local structure. Recall that a double phosphorylated peptide bears two serines, each of which carries one phosphate group cleavable by LPP independently of the other. Under the action of LPP, peptides can be non-phosphorylated (two phosphate groups cleaved) or mono phosphorylated (only one phosphate group cleaved). Non-phosphorylated peptides are however required to trigger phase separation 34. As dephosphorylation proceeds, more RNA chains are incorporated into condensates which concommitantly become more compact. After some maturation period, perhaps once non-phosphorylated peptides are supersaturated, condensate size abruptly increases upon coalescence, until steady state is reached.
Elucidating the nonequilibrium dynamics of biomolecular condensates is essential for understanding the functions and the pathological significance of membraneless organelles in the cell, as well as to design synthetic liquid-like carriers for intracellular delivery of therapeutic mRNA for instance 39. Even though the phase diagram of multicomponent fluid mixtures is inherently complex, particularly in the presence of charged polymers 40, 41, 42, the growth kinetics of our passively-formed condensates obey the yet-simple droplet coarsening law ($\langle R\rangle\sim t^{1/3}$) for binary mixtures. Introducing enzymatic activity delays the onset of the reaction-limited growth, but more importantly, kinetic effects modulate the local structure of condensates. A loosely-packed content may be crucial for preventing dynamical arrest that can occur with the highly charged biomolecules found in the cell nucleus 43. In that sense, enzymatic activity may not only regulate the formation and dissolution of biomolecular condensates, but also the viscoelastic properties that govern nanoscale dynamics 44, 43.
A recent study has reported that kinetics govern the large-scale spatial organization of phase-separated DNA droplets 45. In particular, fast condensation was found to promote the formation of hyperuniform structures. Although our passively-formed condensates grow on a faster timescale, we find no clear evidence of hyperuniformity, neither on scattering curves nor on confocal microscopy images (Fig. LABEL:fig:SI-structure_factor_passive_active). Therefore, the other factors that govern the large-scale spatial organization of phase-separating fluids are yet to be identified.
At last, a number of condensates, notably transcriptional condensates, are immersed in a crowded environment that can be modeled as heterogeneous elastic medium. Recent theoretical advances have uncovered a rich diversity of condensate growth kinetics governed by local environmental stiffness 46. Investigating these effects experimentally using model condensates will be an important direction for future research.
{acknowledgement}
We thank Laetitia Gargowitsch for polyU characterizations, Raphaël Maire for his help on hyperuniformity and Antoine Maury for confocal microscopy observations. We have also benefited from fruitful discussions with Pau Bernado, Christine Doucet and Aurélie Fournet. This work is supported by the Agence Nationale de la Recherche (contract ANR-21-CE30-0001). G.T. acknowledges the ESRF for allocating beamtime. E.R.S is grateful to FAPESP for a visiting research scholarship at Université Paris-Saclay (#23/018385-8).
{suppinfo}
Experimental methods, characterizations of RNA and peptides, supplementary figures with MD simulations, TR-USAXS data and confocal microscopy images.